OpenStudy (anonymous):

Simplify. 1/2-i plz help!

OpenStudy (anonymous):

so its .5-i. i dont think you cant simplify that unless you want to do (1/2)-(sqrt(-1))

OpenStudy (anonymous):

Multiply numerator and denominator by the conjugate of 2-i which is 2+i

OpenStudy (anonymous):

\[\frac{1}{2-i}*\frac{2+i}{2+i}\] \[\frac{2+i}{2+1}\] \[\frac{2+i}{3}\] \[\frac{2}{3}+\frac{1}{3}i\]

OpenStudy (anonymous):

\[(2-i)(2+i)\] \[4+2i-2i-i^2\] \[i^2=-1\] \[4-(-1)\] \[4+1=5\]

OpenStudy (anonymous):

Sorry \[\frac{2}{5}+\frac{1}{5}i\]

OpenStudy (anonymous):

Or is the question: \[\frac{1}{2}-i\] ?